OpenStudy (anonymous):

10.2=x^2*√3/4 solve for x

OpenStudy (anonymous):

see this e.g. ques. x+3/ x-2 + 2 = x+5/x-2 answer.. Multiply each term by the LCD of x-2 thus producing x+3 +2(x-2)=x+5 x+3+2x-4=x+5 3x-1=x+5 2x= 6 x=3 it q=will help u a lot! :)

OpenStudy (anonymous):

10.2 = 3x^2√4 10.2/√4 = 3x^2 10.2/3√4 = x^2 ±√(10.2/3√4) = x

OpenStudy (lgbasallote):

let me get this straight... 10.2 = (x^2)(\[3 sqrt 4) right?

OpenStudy (lgbasallote):

or is it cube root 4?

OpenStudy (anonymous):

it is the formula to find the area of a equalatrial triangle.

OpenStudy (anonymous):

@Rohangrr check your answer again. It comes out with x^2*3√4=54

OpenStudy (anonymous):

Sorry, I see it was an example

OpenStudy (anonymous):

@psujono don't do shanky! it's an EXAMPLE

OpenStudy (anonymous):

\[x=\pm \sqrt{\frac{10.2}{3\sqrt{4}}}\]

OpenStudy (anonymous):

No. wrong again (didn't read question properly)

OpenStudy (anonymous):

basically the 10.2m^2 is the area of an equalatiral trinagle and the original question was the find the lenth of the sides.

OpenStudy (anonymous):

I got the same thing my but text gives a different answer.

OpenStudy (anonymous):

Ah text book answers

OpenStudy (lgbasallote):

did it come out \[x = 2\sqrt{\frac {10.2}{\sqrt{3}}}\] because that would make sense...

OpenStudy (lgbasallote):

wait...i see it can still be simplified... \[x = 2\sqrt{3.4\sqrt{3} }\] right?

OpenStudy (anonymous):

strangely my text gives the answer 4.85m

OpenStudy (lgbasallote):

yuppp if you use calculator..you get 4.85...i used wolfram alpha and this came out http://www.wolframalpha.com/input/?i=2+sqrt+%283.4+sqrt+%283%29+%29

OpenStudy (anonymous):

thanks for guys.